Roman, Bob and Roger buy red and green marbles. Red and green marbles don’t have the same price. Roman pays 70 ¢ for four red marbles and two green marbles. Bob pays 95 ¢ for two red marbles and five green marbles. How much will Roger pay if he buys one red marble and one green marble?

Answer by mananth(16946)

Roman, Bob and Roger buy red and green marbles. Red and green marbles don’t have the same price. Roman pays 70 ¢ for four red marbles and two green marbles. Bob pays 95 ¢ for two red marbles and five green marbles. How much will Roger pay if he buys one red marble and one green marble?

let each red marble cost x
let each green marble cost y
4x+2y=70
2x+5y=95
4x+2y -2(2x+5y)=70-190
4x+2y-4x-10y = -120
-8y= -120
y=-120/-8
y=15 green marble
4x+2*15=70
4x+30=70
4x=40
x=10 red marble
1 red &1 green will cost 25 cents